# Automating Android Build Repair: Bridging the Reasoning-Execution Gap in LLM Agents with Domain-Specific Tools

### v1

@@ 1 Introduction
Android is the largest and most active mobile development platform in the world, with a market share of 73.9% of the global mobile OS market, over 2.2 million apps on the Google Play Store, and over 1.4 million questions tagged "android" on Stack Overflow (Statista, 2024; StatCounter Global Stats, ; Stack Overflow, ). Despite its popularity, a practical challenge remains: applications cannot reliably be built. A comprehensive study of 5,222 apps revealed that only 31.3% could be built successfully out-of-the-box (Liu et al., 2024a). In our dataset, we find the primary source of issues is syntax errors (Table 3(a)). While Continuous Integration and Continuous Deployment (CI/CD) pipelines are a well-established practice for mitigating these errors (Rostami Mazrae et al., 2023; Baitha et al., 2024; Ghaleb et al., 2024; Liu et al., 2022), Liu et al. (2024a) reported only 21.8% out of the 5,222 apps had CI/CD scripts. Furthermore, even with CI/CD, maintaining a stable build requires significant effort. Hassan et al. (2017) found that over 74% of CI/CD configuration changes are dedicated to build fixing and environment updates. This maintenance is necessary to address a complex range of errors, including syntax mistakes, missing resource files, configuration errors, and library unavailability.
Large Language Models (LLMs) and their application in Automated Program Repair (APR) offer a promising alternative (Zhang et al., 2024a; Yin et al., 2024; Xia et al., 2023; Jin et al., 2023). However, the application of LLMs to Android build errors remains an underexplored area. Existing LLM coding benchmarks, like SWE-bench (Jimenez et al., 2023), primarily evaluate models on resolving bugs from GitHub issues, not the transient, environment-specific build failures common in Android development. Furthermore, existing datasets of Android build failures (Liu et al., 2024a; Hassan et al., 2017) capture a repository at a single, broken point in time, making fixes difficult to verify as feasible without a corresponding working build to serve as ground truth.
@@
Second, we propose GradleFixer, an agent-based approach for build error fixing using domain-specific tools. Inspired by prior work showing that specialized tools improve agent performance (Singh et al., 2025; Wang et al., 2023) (details in Appendix A), GradleFixer equips an LLM agent with domain-specific tools, which are simple wrappers for shell commands, to inspect and manipulate the Gradle build environment. Our experiments show that GradleFixer significantly outperforms state-of-the-art baselines across all categories of build failures (Figure 1). We infer from GradleFixer’s success that LLMs have the high-level knowledge needed to fix build errors. The poorer performance of shell-based agents, however, suggests this knowledge fails to transfer to effective low-level execution. This gap explains why other effective agents also rely on specialized toolsets (Wang et al., 2023; Singh et al., 2025), although these studies do not explore this choice as a research question. Our approach, which we term Tool Bridging, addresses this gap by replacing a general-purpose shell with domain-aware abstractions to bridge high-level reasoning to correct low-level execution. We hypothesize this strategy improves performance through two simultaneous mechanisms. First, by providing API-like tools that abstract away complex command syntax, it allows the model to focus on reasoning rather than implementation. Second, it constrains the action space, reducing irrelevant operations. By reframing the task into a format the model uses more reliably (Liu et al., 2024b; Zhang et al., 2024b, 2025), our strategy connects the LLM’s high-level reasoning to effective low-level execution. Our contributions are summarized as follows:
@@ 2.1 Project Curation
AndroidBuildBench is curated from 43 active Android projects on GitHub. These were selected by filtering the top 100 most-starred applications for Java/Kotlin usage, >500 stars, >100 closed pull requests, recent commits, and a Gradle build script. This process yielded 43 open-source projects (Appendix G). A summary of the project statistics is shown in Table 1. We focus on the Gradle build system, as it is used by 95.1% of Android applications according to a recent study (Liu et al., 2024a).
@@ 2.3 Error Categorization
We manually analyzed the 184 test set issues, adopting the five build failure categories proposed by Liu et al. (2024a).
@@
Our analyses are presented in Table 3(a). A notable finding is the absence of NDK errors. We posit this absence results from our curation method, which is anchored to successfully built pull requests. NDK errors often result from build machine environment incompatibilities. Since our method requires a successful build as an endpoint, it naturally filters out states with systemic environment issues. The errors captured are those transiently introduced and fixed within a pull request, a scope more representative of the typical software development life cycle. Furthermore, we find that syntax errors are the most common failure type. This is surprisingly high for human-committed errors, given that modern IDEs typically provide real-time syntax checking. For LLM-generated errors, this result is less surprising, as current models are known to be prone to syntactical mistakes (Tambon et al., 2025).
@@ 3.1 Motivation: Providing the Right Tools
LLMs demonstrate strong proficiency with common shell commands and can often infer how to use novel tools from their descriptions (Patil et al., 2025; Zeng et al., 2025; Team, 2025). However, our preliminary analysis revealed a critical limitation when applying this general capability to a specialized domain like Android build repair. Although an LLM may possess the high-level knowledge to understand a build failure, it struggles to translate this knowledge into the correct sequence of low-level shell commands. For example, correctly synthesizing and sequencing a complex command like ./gradlew assembleDebug along with necessary environment variables is highly error-prone compared to the high-level decision of when to run a build.
We find this gap between high-level reasoning and low-level execution (Tables 4 & 6) likely occurs because while common shell commands (e.g., ls, grep) are prevalent in training data (Liu et al., 2024b; Patil et al., 2024), the multi-step sequences required for domains like Android builds are less common. An agent with a general shell often knows the relevant commands individually but fails to apply them in the correct context, leading to inefficient trial-and-error loops. This phenomenon is analogous to providing a specialist with a curated set of instruments rather than an exhaustive, unorganized toolkit.
Our approach addresses this through Tool Bridging, which replaces general-purpose tools with domain-aware abstractions. These specialized tools translate shell commands into high-level, API-like actions, presenting a format that LLMs handle more reliably and constraining the action space (Patil et al., 2025). This design allows the model to focus its reasoning on debugging rather than command synthesis, bridging the gap between conceptual understanding and correct execution.
@@ 3.2 Execution Environment
All experiments were conducted on a standardized Linux machine (256 CPU cores, 512 GB RAM) installed with the Android SDK. To accommodate the diverse requirements of different projects, we installed the major versions of the Java Development Kit (JDK), including versions 11, 17, 20, 21, 22, and 23. Following the methodology of prior work (Liu et al., 2024a), we build each application using the command ./gradlew assembleDebug --parallel. Building Android applications is computationally intensive, with build times for some projects extending up to 15 minutes on a 256-core machine. We containerized each build process within an isolated Jupyter Notebook (Kluyver et al., 2016). This approach enabled us to run multiple tasks concurrently, significantly improving experimentation throughput. Additionally, to ensure a clean state between runs, we run ./gradlew clean --stop before attempting to build, preventing cached states from affecting the current build.
@@ 3.3 Approaches
We evaluate the effectiveness of an agent-based approach to repairing Android build errors by comparing several configurations designed to isolate the impact of domain-specific tooling. All of our agent configurations use Gemini-2.5-Pro (Comanici et al., 2025) as the core LLM.
As a baseline representing a state-of-the-art coding agent, we use the official, unmodified Gemini-CLI open-source implementation (Google Cloud, ). To allow for controlled experiments, we also developed a validated Python replica of the Gemini-CLI agent (Appendix B). This replica preserves the original agentic loop and prompting (Appendix C) but allows us to easily customize the available toolset. We use this replica for the Gemini-CLI (Read/Write Only) configuration and our proposed method, GradleFixer. We evaluate the following five distinct approaches:
  • Coding-Assistant: A baseline that simulates a user-in-the-loop assistant. The agent is provided with file context and an error log but can only propose code modifications. It cannot execute commands, browse the file system, or run the build. We use the open-source Aider (Aider, ) framework for this configuration.
  • Hierarchical Agent: A two-agent framework where a primary LLM agent reasons about the problem and then delegates code editing tasks to the Coding-Assistant by invoking it as a tool, similar to Ishibashi and Nishimura (2024); Liu et al. (2024c).
@@
  • Gemini-CLI (Read/Write and Shell): The standard open-source Gemini-CLI agent Google Cloud . We use the unmodified open-source version for this baseline. It is given file read/write capabilities and a general-purpose shell tool, allowing it to execute arbitrary commands.
@@ 4.1 Impact of domain-specific tools
Our results show a clear trend of improved performance as tools become more specific. The agent with only the shell tool improves over the baseline. However, performance increases with the more abstract TOOL_B tool and is highest with the most specific tool, TOOL_A. This result is particularly notable because shell is a superset of TOOL_B, which is a superset of TOOL_A. The agent is most effective when given the most constrained, highest-abstraction tool, suggesting the performance gain comes from focus, not just capability. Performance further improved as tools were combined, with our full method achieving the highest resolve rate and demonstrating the value of well-designed shell abstractions. Furthermore, our full method, which excludes the shell, outperforms the combination that includes it (74.0% vs 70.7%), suggesting that a constrained and relevant toolset is more effective than an unrestricted one. These observations align with findings in Wang et al. (2023), where specialized toolsets enable continuous performance improvement while more generic agents plateau.
@@
Tool Bridging addresses this gap by providing domain-aware abstractions. By replacing the general shell with specific tools like TOOL_A, we reframe complex shell interactions as simple, API-like calls. This aligns with the format LLMs are trained to use reliably (Patil et al., 2025) and constrains the action space to relevant operations. This allows the model to leverage its high-level reasoning without failing on low-level command syntax and sequencing. Our results support this hypothesis. As shown in Table 5, performance improves as tools become more specific, suggesting LLMs know which commands are necessary but struggles to sequence them correctly as tools become less specific. GradleFixer achieves superior performance while calling TOOL_C far less frequently (0.6%), suggesting effective targeted tool use rather than exploratory use.
@@
We also include an evaluation using GPT-5-Mini to show that our finding is not limited to a single family of LLMs. While OpenAI provides its own coding agent, Codex OpenAI (2025), it is not open-source, so its exact toolset cannot be reproduced. We instead use the same tools as Gemini-CLI. The results are shown in Table 7.
@@ 4.3 Model Size and Cost-Effectiveness
This principle also supports practical cost-saving strategies like a cascading agent design, where tasks are first attempted by a smaller, cheaper model and only escalated to a larger one upon failure. More broadly, this finding points toward a promising future where smaller, specialized models could be fine-tuned to achieve high performance on domain-specific tasks, potentially exceeding the capabilities of larger generalist models at a much lower cost (Belcak et al., 2025).
@@ 5 Conclusion and Future Work
In conclusion, to address the high frequency of build failures in Android development, we introduced AndroidBuildBench, a benchmark of real-world build errors, and GradleFixer, an LLM agent with domain-specific tools. Our results demonstrate that providing specialized tools is more effective than relying on a general-purpose shell for Android build errors. GradleFixer significantly outperformed state-of-the-art baselines, and we found that a smaller, more cost-effective model with our toolset outperformed a larger model without it. The effectiveness of Tool Bridging also offer a potential explanation for the strong performance of agents in other domains that use special API-like tools (Wang et al., 2023; Singh et al., 2025). Our work provides the empirical evidence for this design pattern, suggesting it is a generalizable approach for building capable LLM agents.
@@
Finally, our explanation for the success of Tool Bridging is a hypothesis supported by strong empirical results. Proving the exact internal reasoning of the language model remains a significant challenge in AI research, and our study does not address the deeper, mechanistic question behind our central finding. For instance, inspired by recent work on the internal mechanisms of LLMs (Lindsey et al., 2025), a deeper investigation could probe whether a dedicated tool like TOOL_A activates a different, more specialized “computational circuit” within the model compared to a generic shell command. Such an analysis would be necessary to connect the behavioral success we observed with its underlying cognitive processes. Therefore, our claim is an inference based on observed performance rather than a direct measurement of the model’s cognitive process. Furthermore, our experimental design deliberately excluded access to commit history to create a focused repair task. This differs from a real world scenario where a developer might use version history to diagnose a problem.
@@ Appendix A Tool Bridging: Relation to Prior Work on Specialized Tools
While we are inspired by prior work that successfully employs specialized tools (Wang et al., 2023; Singh et al., 2025), their focus differs from ours. In prior work, such tools are a foundational component for achieving a broader goal, such as lifelong learning or deep code analysis. In contrast, our work isolates the act of tool replacement itself as the central research question, empirically demonstrating its efficacy.
Distinction from Voyager (Wang et al., 2023). The Voyager agent is an influential paper for lifelong learning within an embodied environment (Minecraft). It operates by building a growing skill library from executable JavaScript functions which call predefined, domain-specific APIs like mineBlock(bot, name). These APIs form the agent’s fundamental action space.
@@
Distinction from Code Researcher (Singh et al., 2025). Code Researcher is a powerful agent designed to patch crashes in complex systems code. It uses specialized tools for code analysis, such as search_code(regex) and search_commits(regex), which are domain-aware for code and history investigation. Similar to Voyager, the paper’s contribution is its deep research methodology, not a comparative analysis of its tools. The research question is not "Is search_commits better than git log in a shell?" but rather "How can an agent use a specialized tool to leverage codebase history to solve complex bugs?". Conversely, our work focuses directly on this comparison. We provide evidence that abstracting complex commands into a constrained, API-like action space is a key strategy for improving agent performance in the Android build domain.

### v2

@@ 1 Introduction
Android is the largest and most active mobile development platform in the world, with a market share of 73.9% of the global mobile OS market, over 2.2 million apps on the Google Play Store, and over 1.4 million questions tagged "android" on Stack Overflow (statista_googleplay_2024; statcounter_os_share; stackoverflow_android_questions). Despite its popularity, a practical challenge remains: applications cannot reliably be built. A comprehensive study of 5,222 apps revealed that only 31.3% could be built successfully out-of-the-box (liu2024understanding). In our dataset, we find the primary source of issues is syntax errors (Table 3(a)). While Continuous Integration and Continuous Deployment (CI/CD) pipelines are a well-established practice for mitigating these errors (rostami2023usage; baitha2024streamlining; ghaleb2024ci; liu2022first), liu2024understanding reported only 21.8% out of the 5,222 apps had CI/CD scripts. Furthermore, even with CI/CD, maintaining a stable build requires significant effort. hassan2017automatic found that over 74% of CI/CD configuration changes are dedicated to build fixing and environment updates. This maintenance is necessary to address a complex range of errors, including syntax mistakes, missing resource files, configuration errors, and library unavailability.
Large Language Models (LLMs) and their application in Automated Program Repair (APR) offer a promising alternative (zhang2024systematic; yin2024thinkrepair; xia2023automated; jin2023inferfix). However, the application of LLMs to Android build errors remains an underexplored area. Existing LLM coding benchmarks, like SWE-bench (jimenez2023swe), primarily evaluate models on resolving bugs from GitHub issues, not the transient, environment-specific build failures common in Android development. Furthermore, existing datasets of Android build failures (liu2024understanding; hassan2017automatic) capture a repository at a single, broken point in time, making fixes difficult to verify as feasible without a corresponding working build to serve as ground truth.
@@
Second, we propose GradleFixer, an agent-based approach for build error fixing using domain-specific tools. Inspired by prior work showing that specialized tools improve agent performance (singh2025code; wang2023voyager) (details in Appendix A), GradleFixer equips an LLM agent with domain-specific tools, which are simple wrappers for shell commands, to inspect and manipulate the Gradle build environment. Our experiments show that GradleFixer significantly outperforms state-of-the-art baselines across all categories of build failures (Figure 1). We infer from GradleFixer’s success that LLMs have the high-level knowledge needed to fix build errors. The poorer performance of shell-based agents, however, suggests this knowledge fails to transfer to effective low-level execution. This gap explains why other effective agents also rely on specialized toolsets (wang2023voyager; singh2025code), although these studies do not explore this choice as a research question. Our approach, which we term Tool Bridging, addresses this gap by replacing a general-purpose shell with domain-aware abstractions to bridge high-level reasoning to correct low-level execution. We hypothesize this strategy improves performance through two simultaneous mechanisms. First, by providing API-like tools that abstract away complex command syntax, it allows the model to focus on reasoning rather than implementation. Second, it constrains the action space, reducing irrelevant operations. By reframing the task into a format the model uses more reliably (liu2024toolace; zhang2024ecoact; zhang2025nemotron), our strategy connects the LLM’s high-level reasoning to effective low-level execution. Our contributions are summarized as follows:
@@ 2.1 Project Curation
AndroidBuildBench is curated from 43 active Android projects on GitHub. These were selected by filtering the top 100 most-starred applications for Java/Kotlin usage, >500 stars, >100 closed pull requests, recent commits, and a Gradle build script. This process yielded 43 open-source projects (Appendix G). A summary of the project statistics is shown in Table 1. We focus on the Gradle build system, as it is used by 95.1% of Android applications according to a recent study (liu2024understanding).
@@ 2.3 Error Categorization
We manually analyzed the 184 test set issues, adopting the five build failure categories proposed by liu2024understanding.
@@
Our analyses are presented in Table 3(a). A notable finding is the absence of NDK errors. We posit this absence results from our curation method, which is anchored to successfully built pull requests. NDK errors often result from build machine environment incompatibilities. Since our method requires a successful build as an endpoint, it naturally filters out states with systemic environment issues. The errors captured are those transiently introduced and fixed within a pull request, a scope more representative of the typical software development life cycle. Furthermore, we find that syntax errors are the most common failure type. This is surprisingly high for human-committed errors, given that modern IDEs typically provide real-time syntax checking. For LLM-generated errors, this result is less surprising, as current models are known to be prone to syntactical mistakes (tambon2025bugs).
@@ 3.1 Motivation: Providing the Right Tools
LLMs demonstrate strong proficiency with common shell commands and can often infer how to use novel tools from their descriptions (patil2025bfcl; zeng2025glm; tbench_2025). However, our preliminary analysis revealed a critical limitation when applying this general capability to a specialized domain like Android build repair. Although an LLM may possess the high-level knowledge to understand a build failure, it struggles to translate this knowledge into the correct sequence of low-level shell commands. For example, correctly synthesizing and sequencing a complex command like ./gradlew assembleDebug along with necessary environment variables is highly error-prone compared to the high-level decision of when to run a build.
We find this gap between high-level reasoning and low-level execution (Tables 4 & 6) likely occurs because while common shell commands (e.g., ls, grep) are prevalent in training data (liu2024toolace; patil2024gorilla), the multi-step sequences required for domains like Android builds are less common. An agent with a general shell often knows the relevant commands individually but fails to apply them in the correct context, leading to inefficient trial-and-error loops. This phenomenon is analogous to providing a specialist with a curated set of instruments rather than an exhaustive, unorganized toolkit.
Our approach addresses this through Tool Bridging, which replaces general-purpose tools with domain-aware abstractions. These specialized tools translate shell commands into high-level, API-like actions, presenting a format that LLMs handle more reliably and constraining the action space (patil2025bfcl). This design allows the model to focus its reasoning on debugging rather than command synthesis, bridging the gap between conceptual understanding and correct execution.
@@ 3.2 Execution Environment
All experiments were conducted on a standardized Linux machine (256 CPU cores, 512 GB RAM) installed with the Android SDK. To accommodate the diverse requirements of different projects, we installed the major versions of the Java Development Kit (JDK), including versions 11, 17, 20, 21, 22, and 23. Following the methodology of prior work (liu2024understanding), we build each application using the command ./gradlew assembleDebug --parallel. Building Android applications is computationally intensive, with build times for some projects extending up to 15 minutes on a 256-core machine. We containerized each build process within an isolated Jupyter Notebook (kluyver2016jupyter). This approach enabled us to run multiple tasks concurrently, significantly improving experimentation throughput. Additionally, to ensure a clean state between runs, we run ./gradlew clean --stop before attempting to build, preventing cached states from affecting the current build.
@@ 3.3 Approaches
We evaluate the effectiveness of an agent-based approach to repairing Android build errors by comparing several configurations designed to isolate the impact of domain-specific tooling. All of our agent configurations use Gemini-2.5-Pro (comanici2025gemini) as the core LLM.
As a baseline representing a state-of-the-art coding agent, we use the official, unmodified Gemini-CLI open-source implementation (google_gemini_cli). To allow for controlled experiments, we also developed a validated Python replica of the Gemini-CLI agent (Appendix B). This replica preserves the original agentic loop and prompting (Appendix C) but allows us to easily customize the available toolset. We use this replica for the Gemini-CLI (Read/Write Only) configuration and our proposed method, GradleFixer. We evaluate the following five distinct approaches:
  • Coding-Assistant: A baseline that simulates a user-in-the-loop assistant. The agent is provided with file context and an error log but can only propose code modifications. It cannot execute commands, browse the file system, or run the build. We use the open-source Aider (aider-website) framework for this configuration.
  • Hierarchical Agent: A two-agent framework where a primary LLM agent reasons about the problem and then delegates code editing tasks to the Coding-Assistant by invoking it as a tool, similar to ishibashi2024self; liu2024towards.
@@
  • Gemini-CLI (Read/Write and Shell): The standard open-source Gemini-CLI agent google_gemini_cli. We use the unmodified open-source version for this baseline. It is given file read/write capabilities and a general-purpose shell tool, allowing it to execute arbitrary commands.
@@ 4.1 Impact of domain-specific tools
Our results show a clear trend of improved performance as tools become more specific. The agent with only the shell tool improves over the baseline. However, performance increases with the more abstract TOOL_B tool and is highest with the most specific tool, TOOL_A. This result is particularly notable because shell is a superset of TOOL_B, which is a superset of TOOL_A. The agent is most effective when given the most constrained, highest-abstraction tool, suggesting the performance gain comes from focus, not just capability. Performance further improved as tools were combined, with our full method achieving the highest resolve rate and demonstrating the value of well-designed shell abstractions. Furthermore, our full method, which excludes the shell, outperforms the combination that includes it (74.0% vs 70.7%), suggesting that a constrained and relevant toolset is more effective than an unrestricted one. These observations align with findings in wang2023voyager, where specialized toolsets enable continuous performance improvement while more generic agents plateau.
@@
Tool Bridging addresses this gap by providing domain-aware abstractions. By replacing the general shell with specific tools like TOOL_A, we reframe complex shell interactions as simple, API-like calls. This aligns with the format LLMs are trained to use reliably (patil2025bfcl) and constrains the action space to relevant operations. This allows the model to leverage its high-level reasoning without failing on low-level command syntax and sequencing. Our results support this hypothesis. As shown in Table 5, performance improves as tools become more specific, suggesting LLMs know which commands are necessary but struggles to sequence them correctly as tools become less specific. GradleFixer achieves superior performance while calling TOOL_C far less frequently (0.6%), suggesting effective targeted tool use rather than exploratory use.
@@
We also include an evaluation using GPT-5-Mini to show that our finding is not limited to a single family of LLMs. While OpenAI provides its own coding agent, Codex openai-codex, it is not open-source, so its exact toolset cannot be reproduced. We instead use the same tools as Gemini-CLI. The results are shown in Table 7.
@@ 4.3 Model Size and Cost-Effectiveness
This principle also supports practical cost-saving strategies like a cascading agent design, where tasks are first attempted by a smaller, cheaper model and only escalated to a larger one upon failure. More broadly, this finding points toward a promising future where smaller, specialized models could be fine-tuned to achieve high performance on domain-specific tasks, potentially exceeding the capabilities of larger generalist models at a much lower cost (belcak2025smalllanguagemodelsfuture).
@@ 5 Conclusion and Future Work
In conclusion, to address the high frequency of build failures in Android development, we introduced AndroidBuildBench, a benchmark of real-world build errors, and GradleFixer, an LLM agent with domain-specific tools. Our results demonstrate that providing specialized tools is more effective than relying on a general-purpose shell for Android build errors. GradleFixer significantly outperformed state-of-the-art baselines, and we found that a smaller, more cost-effective model with our toolset outperformed a larger model without it. The effectiveness of Tool Bridging also offer a potential explanation for the strong performance of agents in other domains that use special API-like tools (wang2023voyager; singh2025code). Our work provides the empirical evidence for this design pattern, suggesting it is a generalizable approach for building capable LLM agents.
@@
Finally, our explanation for the success of Tool Bridging is a hypothesis supported by strong empirical results. Proving the exact internal reasoning of the language model remains a significant challenge in AI research, and our study does not address the deeper, mechanistic question behind our central finding. For instance, inspired by recent work on the internal mechanisms of LLMs (lindsey2025biology), a deeper investigation could probe whether a dedicated tool like TOOL_A activates a different, more specialized “computational circuit” within the model compared to a generic shell command. Such an analysis would be necessary to connect the behavioral success we observed with its underlying cognitive processes. Therefore, our claim is an inference based on observed performance rather than a direct measurement of the model’s cognitive process. Furthermore, our experimental design deliberately excluded access to commit history to create a focused repair task. This differs from a real world scenario where a developer might use version history to diagnose a problem.
@@ Appendix A Tool Bridging: Relation to Prior Work on Specialized Tools
While we are inspired by prior work that successfully employs specialized tools (wang2023voyager; singh2025code), their focus differs from ours. In prior work, such tools are a foundational component for achieving a broader goal, such as lifelong learning or deep code analysis. In contrast, our work isolates the act of tool replacement itself as the central research question, empirically demonstrating its efficacy.
Distinction from Voyager (wang2023voyager). The Voyager agent is an influential paper for lifelong learning within an embodied environment (Minecraft). It operates by building a growing skill library from executable JavaScript functions which call predefined, domain-specific APIs like mineBlock(bot, name). These APIs form the agent’s fundamental action space.
@@
Distinction from Code Researcher (singh2025code). Code Researcher is a powerful agent designed to patch crashes in complex systems code. It uses specialized tools for code analysis, such as search_code(regex) and search_commits(regex), which are domain-aware for code and history investigation. Similar to Voyager, the paper’s contribution is its deep research methodology, not a comparative analysis of its tools. The research question is not "Is search_commits better than git log in a shell?" but rather "How can an agent use a specialized tool to leverage codebase history to solve complex bugs?". Conversely, our work focuses directly on this comparison. We provide evidence that abstracting complex commands into a constrained, API-like action space is a key strategy for improving agent performance in the Android build domain.